During 2018, LCA paid interest of $50,000. At the 2018 fiscal year-end the company's long-term debt was $122,580. The previous year-end (2017), LCA had 156,890 in long-term debt. Calculate the cash flow from/to debtholders for 2018 for the Financial Cash Flows $34,310 -$34,310 O-$17,155 $84,310 -$84,310
O breakeven 1 pts Question 2 At Stage 2 of the decision tree it shows that if a project is successful, the payoff will be $53,000 with a 0.7 chance of occurrence. There is also the 0.3 chance of a -$24,000 payoff. Stage 2 is a year from now. The cost of the project is $25,000 today (Stage 1). The cost of capital is 15 percent. What is the NPV of the project at Stage 1? (Hint: Solve for the expected value of the payoff a year from now, discount the value, and then compare with initial investment). $4,900 -$4,900 $1,000 O -$1,000 -$12,391.30 1 pts
Answer #1

(1) Debt at the end of 2017 = $ 156890 and Debt at the end of 2018 = $ 122580

A reduction in debt implies that a portion of it has been paid off during 2018 by LCA to debtholders.

Cash Flow to Debtholders = 156890 - 122580 = $ 34310 and Interest Payment = $ 50000 (cash flow to debtholders)

Therefore, total cash flow to debtholders = 50000 + 34310 = $ 84310

Hence, the correct option is (d)

NOTE: Please raise separate queries for the solution to the second unrelated question, as one query is restricted to the solution of only one complete question with up to four sub-parts.